OpenStudy (anonymous):

Need help please(last question I have): Find the indefinite integral of: sin^(3)(18x)*cos^(9)(18x) dx

OpenStudy (anonymous):

Do you use this rule: Sin A Cos B = 1/2[sin(A-B) + sin (A+B)]?

OpenStudy (turingtest):

no, strip out a sin(18)x

OpenStudy (turingtest):

sin(18x) I mean...

OpenStudy (turingtest):

sin^(3)(18x)*cos^(9)(18x) dx=sin(18x)*sin^(2)(18x)*cos^(9)(18x) dx

OpenStudy (turingtest):

now note that sin^2x=1-cos^2x

OpenStudy (anonymous):

okay I see what you did

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ya it's not so hard ..i will tell u how to solve this types of problem

OpenStudy (turingtest):

now it's all gonna be cos after simplifying, each multiplied by sin easily done from here with a couple simple u-subs

OpenStudy (anonymous):

so you'll have: sin(18x) *(1-cos^(2)x) * cos^(9)(18x)? dx

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

simplify

OpenStudy (anonymous):

what do you do next? Im not so good at this :(

OpenStudy (turingtest):

distribute and what do you get?

OpenStudy (anonymous):

trying to find my notes for problems like these

OpenStudy (anonymous):

sin(18x) - cos^(2)xsin(18x)

OpenStudy (anonymous):

and the cos^(9)(18x)

OpenStudy (anonymous):

after that, do u do a U substitution of sin 18x?

OpenStudy (turingtest):

write out what you get after distribution completely so I can see where you are

OpenStudy (anonymous):

sin(18x) - cos^(2)(x)sin(18x)*cos^(9)(18x) dx

OpenStudy (turingtest):

\[\sin(18x)(1-\cos^2(18x))(\cos^9(18x))\]\[=\sin(18x)(\cos^2(18x)-\cos^{11}(18x))\]

OpenStudy (anonymous):

okay I see how you got that

OpenStudy (turingtest):

notice that sin(n) is almost the derivative of cos(n)

OpenStudy (anonymous):

yea

OpenStudy (turingtest):

you have to be a little careful here, there is a chain rule element to this u-sub

OpenStudy (anonymous):

yeah, you get cos(8x)(8) right?

OpenStudy (anonymous):

i mean 18x

OpenStudy (anonymous):

why did I write down 8 on my paper -.-. It becomes: cos(18x)(18)

OpenStudy (anonymous):

then du/18 = cos(18x)dx right?

OpenStudy (turingtest):

why do you want to make u=sin(18x) ? how are you going to represent the cos^2 and such in the interval

OpenStudy (turingtest):

?

OpenStudy (turingtest):

integral*

OpenStudy (anonymous):

hmm since du/18 = cos(18x)dx, cant u say (du/18)^(11) is the same as cos^(11)(18x)?

OpenStudy (turingtest):

yeah, and what the heck is the integral\[\int\frac{du^{11}}{18}\]gonna be ?

OpenStudy (turingtest):

what does \(du^n\) mean? nothing, in this context, it's undeined

OpenStudy (anonymous):

what do you suggest U to be?

OpenStudy (turingtest):

you should always consider what's going to happen when you integrate (fx))^n in these cases you should a,ost always try to choose the function that is being raised to a power when it is multiplied by a constant multiple of its derivative

OpenStudy (turingtest):

almost*

OpenStudy (anonymous):

so in this case, it would be what Chlorophyll suggested

OpenStudy (anonymous):

I mean u = cos(18x) -> du = ?

OpenStudy (turingtest):

no, it would be the function inside the power

OpenStudy (turingtest):

I thought you knew that @Chlorophyll you beat me to correcting you ;)

OpenStudy (anonymous):

oh ok, so cos(18x) = -sin(18x)(18)

OpenStudy (turingtest):

you should always write the dx, that illustrates exactly what you are subbing

OpenStudy (anonymous):

yeah sorry

OpenStudy (turingtest):

\[u=\cos(18x)\implies du=-18\sin(18x)dx\]

OpenStudy (anonymous):

u = cos(18x) du = -sin(18x)(18) dx du / 18 = -sin(18x)

OpenStudy (anonymous):

and dx at the end*

OpenStudy (turingtest):

you ought to put that - sign on the other side as well we are subbing for sin(18x)dx, not -sin(18x)dx

OpenStudy (anonymous):

true, -du/18 = sin(18x) dx

OpenStudy (turingtest):

now sub that into what we had way back up there

OpenStudy (anonymous):

(-du/18) (u^(2) - u^(11))

OpenStudy (turingtest):

yep :)

OpenStudy (turingtest):

wait we missed a 9

OpenStudy (anonymous):

( 1 - u² ) u^9

OpenStudy (turingtest):

I sensed Chlorophyll correcting me :)

OpenStudy (anonymous):

xD what do you mean a 9? where?>

OpenStudy (anonymous):

@TuringTest Sorry !

OpenStudy (turingtest):

oh I just distributed wrong being the lazy slob that I am

OpenStudy (anonymous):

ah I see

OpenStudy (turingtest):

oh by all means please, it's great to have more than one pair of eyes on a problem @Chlorophyll

OpenStudy (anonymous):

so how did we get (1-cos(18x))cos^(9)(18x)?

OpenStudy (anonymous):

@TuringTest I'm too too lazy, that's why I'm sort of shadow better person :)

OpenStudy (turingtest):

(1-cos^2(18x))cos^(9)(18x)

OpenStudy (anonymous):

yeah I forgot the ^2

OpenStudy (turingtest):

sin^3x*cos^9x=sinx*sin^2x*cos^9x=sinx*(1-cos^2x)*cos^9x

OpenStudy (turingtest):

strip out a sin to get an even power that we can convert to cosine is the overarching theme of these problems I'd say

OpenStudy (anonymous):

hmm I see how you got what you have now, but what do you do now? -du/18 is sin(18x)dx

OpenStudy (anonymous):

Do we do a u substitution for Cos(x)?

OpenStudy (anonymous):

sinx*(1-cos^2x)*cos^9x Pretend X = 18x Let u = cos 18x

OpenStudy (anonymous):

-> - du/18 = sin 18x => -1/18 ∫ ( 1 - u² ) u^9 du

OpenStudy (anonymous):

so then u get: -1/18[(u-(u^(3) / 3)) (u^(10/10)], correct?

OpenStudy (anonymous):

Note: ( 1 - u² ) u^9 = u^9 - u^11

OpenStudy (anonymous):

yeah so u^(-2)

OpenStudy (anonymous):

No, you need to do integral one by one term!

OpenStudy (anonymous):

ahhh nvm, i just confused myself. Yeah.

OpenStudy (anonymous):

so now we have: 1/18 integral of (u^(9) - u^(11))

OpenStudy (anonymous):

taking the integral, you get: 1/18 [u^(10)/10 - u^(12) / 12 ] right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

= u^(10) / 180 - u^(12) / 216

OpenStudy (anonymous):

Now, put cos18x back!

OpenStudy (anonymous):

= cos^(10)(18x) / 180 - cos^(12)(18x) / 216

OpenStudy (anonymous):

we also forgot a +C

OpenStudy (anonymous):

Watch out the negative sign ( - 1/18)

OpenStudy (anonymous):

Oops.

OpenStudy (anonymous):

okay so: -cos^(10)(18x) / 180 + cos^(12)(18x) / 216

OpenStudy (anonymous):

+ C xD, forgot it again

OpenStudy (anonymous):

Fully correct now :)

OpenStudy (anonymous):

Got it! -((cos^(10)(18x))/180)+ cos^(12)(18x) / 216 is the correct answer. Thank you so much!!

OpenStudy (anonymous):

I'm glad the team work works out :)

OpenStudy (anonymous):

ahaha :)